# Towards the understanding of heavy quarks hadronization: from leptonic to heavy-ion collisions

### v2

@@ I Introduction
The hadrons containing a charm or beauty quark (heavy-flavour hadrons), are always formed in the hadronization of a charm or beauty quark produced in initial hard-scatterings prior to hadronization, in which contrary to the soft interactions occurring at hadronization time, large $Q^{2}$ ($Q^{2}>(2m_{\mathrm{c,b}})^{2}$) are possible. The hadronization process determines the relative abundances of the different heavy-flavour hadron species and the hadron kinematics, but contrary to what occurs for light-flavour hadrons, the hadronization process itself generally does not influence the total heavy-flavour hadron production which instead matches the production rate of the parent heavy quarks. The large values of the charm and beauty quark masses ($m_{b,c}\,>>\,\Lambda_{QCD}$) allow to employ perturbative QCD (pQCD) calculations of heavy-quark production down to zero momentum. For these reasons, heavy-flavour hadrons can be regarded as “calibrated” perturbative probes of the hadronization, which can provide unique information on this fundamental process, complementary to what can be inferred from studies of light-flavour production and of the event properties.
@@
The transition of a heavy quark into a given hadron species requires the modelling of hadronization and of the soft non-perturbative processes involved. The simplest effective description of this transition, adopted in QCD calculations based on a factorization approach like FONLL [3, 4] and GM-VFNS [5, 6, 7, 8], is to use fragmentation functions (FFunc). In the factorization approach, the transverse-momentum ($p_{\mathrm{T}}$) and rapidity ($y$) differential heavy-flavour hadron production cross sections in proton–proton (pp) are computed as the convolution of three terms: (i) the parton distribution functions (PDFs) of the incoming (anti)protons, (ii) the partonic cross section, calculated as a perturbative series in powers of the strong coupling constant $\alpha_{\mathrm{s}}$, and (iii) the fragmentation functions, which parameterize the probability that a heavy-quark gives rise to a hadron of a given species carrying a given fraction ($z_{\mathrm{Q}}$) of the parent quark momentum. Thus, FFunc allow to map the quark momentum-differential cross section into the final hadron momentum-differential cross section. The FFunc used in these pQCD calculations have been parameterized for decades from measurements performed in $\mathrm{e^{+}e^{-}}$ or ep collisions [9]. Charm and bottom quark production in $\mathrm{e^{+}e^{-}}$ annihilation offers the cleanest environment for constraining heavy-quark FFunc. However, the usage of FFunc parameterized on $\mathrm{e^{+}e^{-}}$ or ep data in pp collisions requires the assumption that FFunc are “universal”, independent of the colliding system. Under the universality assumption, once determined in a given collision system, like $\mathrm{e^{+}e^{-}}$, FFunc can be applied to any reaction. An even more basic assumption, is the universality of fragmentation fractions (FF), which represent the probability that a heavy-quark produces a given heavy-flavour hadron species independently of $z_{\mathrm{Q}}$, i.e. the $z_{\mathrm{Q}}$ integral of the FFunc. The modelling of hadronization in hadronic collisions with universal FFunc and FF relies on the hypothesis that, at large extent, the hadronization of a charm or beauty quark is not sensitive to the hard quark-production process and to the surrounding partonic environment in which hadronization occurs. In $\mathrm{e^{+}e^{-}}$ collisions the FF can be directly measured from the cross sections of the observed hadron species without the need of measuring all existing species, given that the heavy-quark production cross section can be calculated with high precision. Though practical limitations (e.g. kinematic reach of the measurements) may require extrapolation, this in principle allows the important check that the sum of FF of all observed weakly-decaying charm hadrons equals unity, thus the verification that all weakly decaying states are known. Though the universality assumptions are not a rigorous prediction of QCD and can be considered reasonable only within certain (quite vaguely defined) kinematic conditions, validating them has been an important goal of experimental measurements of heavy-flavour hadron production since the formulation of QCD and of the parton model. Hadron-to-hadron production cross section ratios are particularly sensitive to FFunc, FF, and in general, to heavy-flavour hadronization mechanisms, because the heavy-quark production cross-section, which depends on PDF and the hard-scattering process, substantially cancel in the ratios. The production of specific heavy-flavour hadron species has been measured in different energy regimes and “hadronic environments”: in $\mathrm{e^{+}e^{-}}$ collisions at LEP and B factories, in ep collisions, in proton-(anti)proton (pp/$\mathrm{\bar{}}{p}$) collisions, and in heavy-ion collisions (nucleus–nucleus, AA, or proton–nucleus, p–A).Before the Large Hadron Collider (LHC) era, the universality assumption had been tested mainly by comparing cross-section ratios of different D and B meson species in $\mathrm{e^{+}e^{-}}$, $\mathrm{e^{\pm}p}$, and pp ($\mathrm{p\overline{p}}$) collisions, while data on baryon-to-meson cross-section ratios in hadronic collisions were not available or precise, and basically limited to $\Lambda_{\mathrm{c}}^{+}$ and $\Lambda_{\mathrm{b}}^{0}$. The $\Lambda_{\mathrm{c}}^{+}/\mathrm{D}$ ratios measured in $\mathrm{e^{\pm}p}$ and $\mathrm{e^{+}e^{-}}$ collisions were consistent, thus supporting universality. Hadronization models developed in event generators, like the Lund string fragmentation model in pythia [10, 11] or the cluster model in HERWIG [12], capable to reproduce several feature of $\mathrm{e^{+}e^{-}}$ data and pp data, were not expecting, in their standard implementation, significant differences in the hadron-species ratios from $\mathrm{e^{+}e^{-}}$ to pp collisions, also supporting the hypothesis that modifications to the hadronization process could be small. On the contrary, the recent heavy-flavour baryon measurements in hadronic collisions at LHC energies indicated a significant difference in the fragmentation fractions of charm and beauty quarks into charm and beauty baryons compared to those measured in $\mathrm{e^{+}e^{-}}$ and ep collisions. In the beauty sector, the CDF experiment had found a difference in the $\Lambda_{\mathrm{b}}^{0}/B$ ratio in $\mathrm{p\overline{p}}$ collisions at Tevatron with respect to LEP values [13]: at that time, however, the result was related to the different jet momenta involved in the two measurements (much higher at LEP) but not traced back to a breakdown of factorization, a modification of the hadronization process, or some other cause. The recent heavy-flavour baryon measurements in hadronic collisions challenge the assumption that heavy quark hadronization is a universal process across different colliding systems. The breaking of the universality can be related to higher-twist effects or other effects that make heavy-quark hadronization dependent on the underlying event.
@@
In heavy-ion collisions, hadronization of heavy quarks (as well as that of light partons) can occur in a rather different way. In these collisions a colour-deconfined state of QCD matter is formed, the quark-gluon plasma (QGP) [14]. The heavy-quark hadronization process happens in a hot and dense environment with a large abundance of light coloured thermal particles nearby with which the heavy quark can recombine, neutralizing its colour charge. Accounting for this recombination process, very often modelled as a 2 → 1 (or 3 → 1 in the case of baryon formation) coalescence is a fundamental hadronization mechanism necessary to reproduce several experimental observations, like changes in baryon-to-meson ratios and the collective motion of final hadrons, which is associated with the expansion of the system, both in the light and the heavy-flavour sector. In heavy-ion collisions also other observables are sensitive to the hadronization processes, like the nuclear modification factor and the anisotropic flow coefficients. The nuclear modification factor ($R_{\mathrm{AA}}$) is defined as the ratio of the yield measured in AA collisions to the pp cross section scaled by the average nuclear overlap function $\langle T_{\mathrm{AA}}\rangle$ [15], a quantity proportional to the number of binary nucleon–nucleon collisions, for the considered centrality interval. In the absence of nuclear effects, $R_{\mathrm{AA}}=1$ is expected. The anisotropic flow coefficients are the Fourier coefficients ($v_{\mathrm{n}}$) of the distribution of the heavy-flavour hadron azimuthal angle ($\phi$) with respect to the reaction plane defined by the nuclei impact-parameter vector, which is the vector connecting the centres of the colliding nuclei, and the beam direction. The coefficient of the second-order harmonic ($v_{\mathrm{2}}$) is called elliptic flow. It encodes the efficiency by which the hydrodynamic evolution of the medium converts to momentum space the initial spatial azimuthal anisotropy of the fireball (and of the produced partons), which, in non-central collisions, principally arises from the non-zero eccentricity of the nuclei overlap region. The initial asymmetry gives rise to azimuthally anisotropic pressure gradients which drives the hydrodynamic expansion of the system.
@@ II Models of hadronization in high-energy collisions
7: In some SHM approaches, but not in those discussed within this review for open heavy flavor.
@@ II.1.1 String model – PYTHIA
Once a flavour and a $p_{\perp{q}}$ for the next string break in the iterative process has been selected, the next step is to select a hadron type compatible with the flavour of the current endpoint and that produced in the breakup. Since spin-1 mesons are heavier than spin-0 ones in QCD, the observed ratios between vectors and pseudoscalars are generally smaller than 3 (and equivalently for spin-3/2 baryons compared to spin-1/2 ones), however the baseline string model does not predict what these factors should be [^1] We note the recent work on making the model more predictive in this regard [23].; instead, these are determined from data, via tuning [^2] In the context of heavy quarks, we note that the default PYTHIA Monash tune has a known issue with the $D^{*}/D$ ratio being too low. This can be improved by increasing the parameter StringFlav:mesonCvector from 0.88 to 1.5, which we recommend for charm studies..
@@ II.2.3 Statistical hadronization with additional baryon resonances
The Statistical Hadronization Model (SHM) assumes that hadrons are produced according to the density of thermal yields $N_{h}^{th}(T)$. In ultra-relativistic heavy-ion collisions, the SHM successfully reproduces the yields of meson and baryon ground states with one chemical freeze-out temperature of $T_{cf}\simeq 156$ MeV [40], which represents the temperature of the medium at which the different particle abundances get fixed. Its extension to charm (and beauty) hadrons needs to consider that the charm quarks are produced in the initial stage of the collisions with a multiplicity that is quite larger than the one expected from thermal charm production at $T_{cf}$. To this purpose, a charm-quark fugacity factor $g_{\mathrm{c}}$, which represents the ratio of total charm yield to the thermal charm one, is included in the model to take into account this excess. A charm balance equation is needed to guarantee the conservation of charm quarks for a given value of $g_{c}$, while allowing the relative abundances of the different open and hidden charm hadrons to follow the statistical thermal ones, it can be written as:
@@
where $N_{\mathrm{c}\bar{c}}$ in the total charm yield in the rapidity range considered, $n_{i,j,k}^{{\mathrm{t}h}}$ are the (grand-canonical) thermal yields for open charm with one charm quark ($h_{oc,1}$) and two charm quarks ($h_{oc,2}$) and for hidden charm ($h_{hc}$). In principle, Eq. 5 should include the hadrons with three charm quarks but, in the current implementations, they are discarded under the reasonable assumption that their contribution to the sum is negligible. Fugacity factors in elementary collisions are introduced to account for the non-equilibrium yield of quarks heavier than $u$ and $d$, thus for strangeness as well as for charm. However, due to the dominance of hadrons with one charm quark this does not affect baryon-to-meson yield ratios of single charm hadrons, because both baryons and mesons contain one charm. On the other hand, one can also notice that doubling $N_{\mathrm{c}\bar{c}}$ implies doubling $g_{\mathrm{c}}$, with a consequent enhancement by a factor of 2 of single-charm hadrons and a factor of 4 increase of double-charm hadrons. Such a scaling with $N_{\mathrm{c}\bar{c}}$ is expected also in the coalescence model described in the next paragraph.
@@ II.2.4 Hadronization via coalescence
with $g_{H}$ indicating the statistical factor giving the probability that two (or three) random quarks have the right quantum numbers to match the quantum number of the hadron (flavour, spin, isospin) in a colourless combination. For D mesons the statistical factor $g_{\mathrm{D}}=1/36$ gives the probability that two random quarks have the right colour, spin, and isospin to match the quantum number of the considered mesons. For $\Lambda_{\mathrm{c}}^{+}$ the statistical factor is $g_{\Lambda_{\mathrm{c}}^{+}}=1/108$. $N_{\mathrm{q}}$ is the number of quarks that form the hadron, $N_{\mathrm{q}}=2$ for mesons and $N_{\mathrm{q}}=3$ for baryons. The $d\sigma_{i}$ denotes an element of a space-like hypersurface, while $f_{{\mathrm{q}}_{i}}$ are the quark (anti-quark) phase-space distribution functions for $i^{\mathrm{th}}$ quark (anti-quark) and the coalescence function ${\cal{C}}_{H}=C^{N_{\mathrm{q}}-1}f_{H}(x_{1}...x_{N_{{\mathrm{q}}}},p_{1}...p_{N_{{\mathrm{q}}}})$ is directly proportional to the Wigner function which describes the spatial and momentum distribution of quarks in a hadron, $f_{H}$. The constant $C^{N_{\mathrm{q}}-1}$ is an important and specific aspect of the Catania model, not present in the standard coalescence approach. It is this constant that allows normalizing the coalescence probability for charm quarks such that the total coalescence probability (i.e. summing up all the mesons and baryons channel considered) in the limit $p_{\mathrm{T}}\rightarrow 0$ is $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$. Such a constant $C$ can be determined only by calculating all the charmed mesons and baryons. In the first work [37] it was evaluated considering only D, $\mathrm{D^{*+}}$, $\mathrm{D_{s}^{+}}$, $\Lambda_{\mathrm{c}}^{+}$, and $\Sigma_{\mathrm{c}}^{0,+,++}$ hadrons. However, more recently the production of $\Xi_{\mathrm{c}}^{+,0}$ and $\Omega_{\mathrm{c}}^{0}$ was observed not to be negligible [36]. Having fixed such a normalization, the probability of a charm quark to hadronize via coalescence or fragmentation depends as in the standard approach on its transverse momentum according to Eq. 6. The coalescence probability is high for low quark momenta and quickly decreases with increasing transverse momentum, meaning that most of charm quarks hadronize via coalescence at low $p_{\mathrm{T}}$ whereas, in the high momentum region, fragmentation is the dominant contribution. Imposing the condition $P_{coal}(p\rightarrow 0)=1$ is a key aspect underlying the prediction of quite large values of heavy baryons: $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}\sim O(1)$, $\Xi_{\mathrm{c}}^{0}/\mathrm{D^{0}}\sim 0.2$ and $\Omega_{\mathrm{c}}^{0}/\mathrm{D^{0}}\sim 0.1$ [36]. The physical motivation is related to the fact that in the standard coalescence approach there is no dynamics related to confinement. The renormalization constant $C$ appears linearly for mesons and quadratic for baryons. This is physically motivated at least by two considerations. First, enforcing $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ implies that there is an enhancement of the recombining probability (with respect to the one from a simple phase space overlap of the wave functions) that can be associated to the dynamics of the confining interaction that increases the coalescence probability (here accounted for just by a renormalization of a coupling constant). Such an effect has to be squared for baryons for which two relative distances among the quarks have to be defined. Second, one has to consider that the density of the medium at which the hadrons are formed is not known, and being the phase transition a cross-over it is likely that there is a density interval over which it occurs. Therefore the hadrons yields and ratios in a standard coalescence approach (i.e. without imposing $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$) would significantly depend on the specific value of the density chosen. In the Catania approach the probability of forming a hadron scales linearly with the density for mesons and quadratically for baryons. If one chooses a lower value density value, implying a reduced standard coalescence probability, the constraint $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ increases the value of the coefficient $C^{N_{q}-1}$, making the yields and the baryon-over-meson ratios quite stable and nearly independent of the specific value of the density of the background fireball employed to perform the calculation.
@@
with $x_{ri}$ and $p_{ri}$ relative coordinates in space and momentum. We notice that an important aspect of a proper implementation of the coalescence approach is that, respecting the basic feature of quantum mechanics, the width $\sigma_{r}$ of the wave function in r-space constraints also the Wigner function width in p-space; for a Gaussian wave function for each relative coordinate we have $\sigma_{ri}\cdot\sigma_{pi}=1$ and $\sigma_{p}$ cannot be used as a further independent parameter. It can be noticed that the coalescence process, and the related formula, can be factorized as a diquark recombination followed by a diquark-quark coalescence if the diquark binding energy is vanishing.
It is interesting to discuss the dependence of the coalescence results on the specific value of the wave-function width, which is the main parameter of the model. This can be easily evaluated analytically considering a non-relativistic reduction, which, even if not quantitatively exact, allows having a semi-quantitative estimate of the main aspect involved in the coalescence process. Assuming a non-relativistic boost invariant thermal distribution function homogeneous in $r$-space, discarding the radial flow, and assuming a Gaussian wave function, it can be shown that the coalescence yield for mesons and baryons scales as:

### v3

@@ I Introduction
The hadrons containing a charm or beauty quark (heavy-flavour hadrons), are always formed in the hadronization of a charm or beauty quark produced prior to hadronization, in initial hard-scatterings or in gluon splittings in parton showers. In these processes, contrary to the soft interactions occurring at hadronization time, large $Q^{2}$ ($Q^{2}>(2m_{\mathrm{c,b}})^{2}$) are possible. The hadronization process determines the relative abundances of the different heavy-flavour hadron species and the hadron kinematics, but, contrary to what occurs for light-flavour hadrons, the hadronization process itself generally does not influence the total heavy-flavour hadron production which instead matches the production rate of the parent heavy quarks. The large values of the charm and beauty quark masses ($m_{b,c}\,>>\,\Lambda_{QCD}$) allow to employ perturbative QCD (pQCD) calculations of heavy-quark production down to zero momentum. For these reasons, heavy-flavour hadrons can be regarded as “calibrated” perturbative probes of the hadronization, which can provide unique information on this fundamental process, complementary to what can be inferred from studies of light-flavour production and of the event properties.
@@
The transition of a heavy quark into a given hadron species requires the modelling of hadronization and of the soft non-perturbative processes involved. The simplest effective description of this transition, adopted in QCD calculations based on a factorization approach like FONLL [3, 4] and GM-VFNS [5, 6, 7, 8], is to use fragmentation functions (FFunc). In the factorization approach, the transverse-momentum ($p_{\mathrm{T}}$) and rapidity ($y$) differential heavy-flavour hadron production cross sections in proton–proton (pp) are computed as the convolution of three terms: (i) the parton distribution functions (PDFs) of the incoming (anti)protons, (ii) the partonic cross section, calculated as a perturbative series in powers of the strong coupling constant $\alpha_{\mathrm{s}}$, and (iii) the fragmentation functions, which parameterize the probability that a heavy-quark gives rise to a hadron of a given species carrying a given fraction ($z_{\mathrm{Q}}$) of the parent quark momentum. Thus, FFunc allow to map the quark momentum-differential cross section into the final hadron momentum-differential cross section. The FFunc used in these pQCD calculations have been parameterized for decades from measurements performed in $\mathrm{e^{+}e^{-}}$ or ep collisions [9]. Charm and bottom quark production in $\mathrm{e^{+}e^{-}}$ annihilation offers the cleanest environment for constraining heavy-quark FFunc. However, the usage of FFunc parameterized on $\mathrm{e^{+}e^{-}}$ or ep data in pp collisions requires the assumption that FFunc are “universal”, independent of the colliding system. Under the universality assumption, once determined in a given collision system, like $\mathrm{e^{+}e^{-}}$, FFunc can be applied to any reaction. An even more basic assumption, is the universality of fragmentation fractions (FF), which represent the probability that a heavy-quark produces a given heavy-flavour hadron species independently of $z_{\mathrm{Q}}$, i.e. the $z_{\mathrm{Q}}$ integral of the FFunc. The modelling of hadronization in hadronic collisions with universal FFunc and FF relies on the hypothesis that, at large extent, the hadronization of a charm or beauty quark is not sensitive to the hard quark-production process and to the surrounding partonic environment in which hadronization occurs. In $\mathrm{e^{+}e^{-}}$ collisions the FF can be directly measured from the cross sections of the observed hadron species without the need of measuring all existing species, given that the heavy-quark production cross section can be calculated with high precision. Though practical limitations (e.g. kinematic reach of the measurements) may require extrapolation, this in principle allows the important check that the sum of FF of all observed weakly-decaying charm or beauty hadrons equals unity, thus the verification that all weakly decaying states are known. Though the universality assumptions are not a rigorous prediction of QCD and can be considered reasonable only within certain (quite vaguely defined) kinematic conditions, validating them has been an important goal of experimental measurements of heavy-flavour hadron production since the formulation of QCD and of the parton model. Hadron-to-hadron production cross section ratios are particularly sensitive to FFunc, FF, and in general, to heavy-flavour hadronization mechanisms, because the heavy-quark production cross-section, which depends on PDF and the hard-scattering process, substantially cancel in the ratios. The production of specific heavy-flavour hadron species has been measured in different energy regimes and “hadronic environments”: in $\mathrm{e^{+}e^{-}}$ collisions at LEP and B factories, in ep collisions, in proton-(anti)proton (pp/$\mathrm{\bar{}}{p}$) collisions, and in heavy-ion collisions (nucleus–nucleus, AA, or proton–nucleus, p–A).Before the Large Hadron Collider (LHC) era, the universality assumption had been tested mainly by comparing cross-section ratios of different D and B meson species in $\mathrm{e^{+}e^{-}}$, $\mathrm{e^{\pm}p}$, and pp ($\mathrm{p\overline{p}}$) collisions, while data on baryon-to-meson cross-section ratios in hadronic collisions were not available or precise, and basically limited to $\Lambda_{\mathrm{c}}^{+}$ and $\Lambda_{\mathrm{b}}^{0}$. The $\Lambda_{\mathrm{c}}^{+}/\mathrm{D}$ ratios measured in $\mathrm{e^{\pm}p}$ and $\mathrm{e^{+}e^{-}}$ collisions were consistent, thus supporting universality. Hadronization models developed in event generators, like the Lund string fragmentation model in pythia [10, 11] or the cluster model in HERWIG [12], capable to reproduce several feature of $\mathrm{e^{+}e^{-}}$ data and pp data, were not expecting, in their standard implementation, significant differences in the hadron-species ratios from $\mathrm{e^{+}e^{-}}$ to pp collisions, also supporting the hypothesis that modifications to the hadronization process could be small. On the contrary, the recent heavy-flavour baryon measurements in hadronic collisions at LHC energies [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] indicated a significant difference in the fragmentation fractions of charm and beauty quarks into charm and beauty baryons compared to those measured in $\mathrm{e^{+}e^{-}}$ and ep collisions. In the beauty sector, the CDF experiment had found a difference in the $\Lambda_{\mathrm{b}}^{0}/B$ ratio in $\mathrm{p\overline{p}}$ collisions at Tevatron with respect to LEP values [29]: at that time, however, the result was related to the different jet momenta involved in the two measurements (much higher at LEP) but not traced back to a breakdown of factorization, a modification of the hadronization process, or some other cause. The LHC measurements challenge the assumption that heavy-quark hadronization is a universal process across different colliding systems. The breaking of the universality can be related to effects that make heavy-quark hadronization dependent on the underlying event.
@@
In heavy-ion collisions, hadronization of heavy quarks (as well as that of light partons) can occur in a rather different way. In these collisions a colour-deconfined state of QCD matter is formed, the quark-gluon plasma (QGP) [30, 31, 32]. The heavy-quark hadronization process happens in a hot and dense environment with a large abundance of light coloured thermal particles nearby with which the heavy quark can recombine, neutralizing its colour charge. The recombination process, frequently modeled as a 2 → 1 coalescence (or 3 → 1 in the case of baryon formation), is a fundamental hadronization mechanism crucial for reproducing several experimental observations. These include shifts in baryon-to-meson ratios and the collective motion of final-state hadrons, linked to the system’s expansion, both in the light and heavy-flavour sectors. In heavy-ion collisions also other observables are sensitive to the hadronization processes, like the nuclear modification factor and the anisotropic flow coefficients. The nuclear modification factor ($R_{\mathrm{AA}}$) is defined as the ratio of the yield measured in AA collisions to the pp cross section scaled by the average nuclear overlap function $\langle T_{\mathrm{AA}}\rangle$ [33], a quantity proportional to the number of binary nucleon–nucleon collisions, for the considered centrality interval. In the absence of nuclear effects, $R_{\mathrm{AA}}=1$ is expected. The anisotropic flow coefficients are the Fourier coefficients ($v_{\mathrm{n}}$) of the distribution of the heavy-flavour hadron azimuthal angle ($\phi$) with respect to the reaction plane defined by the nuclei impact-parameter vector, which is the vector connecting the centres of the colliding nuclei, and the beam direction. The coefficient of the second-order harmonic ($v_{\mathrm{2}}$) is called elliptic flow. It encodes the efficiency by which the hydrodynamic evolution of the medium converts to momentum space the initial spatial azimuthal anisotropy of the fireball (and of the produced partons), which, in non-central collisions, principally arises from the non-zero eccentricity of the nuclei overlap region. The initial asymmetry gives rise to azimuthally anisotropic pressure gradients which drives the hydrodynamic expansion of the system.
@@ II Models of hadronization in high-energy collisions
7: In some SHM approaches, but not in those discussed within this review for open heavy flavour.
@@ II.1.1 String model – PYTHIA
Once a flavour and a $p_{\perp{q}}$ for the next string break in the iterative process has been selected, the next step is to select a hadron type compatible with the flavour of the current endpoint and that produced in the breakup. Since spin-1 mesons are heavier than spin-0 ones in QCD, the observed ratios between vectors and pseudoscalars are generally smaller than 3 (and equivalently for spin-3/2 baryons compared to spin-1/2 ones), however the baseline string model does not predict what these factors should be [^1] We note the recent work on making the model more predictive in this regard [41].; instead, these are determined from data, via tuning [^2] In the context of heavy quarks, we note that the default PYTHIA Monash tune has a known issue with the $\mathrm{D^{*+}}/\mathrm{D^{0}}$ ratio being too low. This can be improved by increasing the parameter StringFlav:mesonCvector from 0.88 to 1.5, which we recommend for charm studies..
@@ II.2.3 Statistical hadronization with additional baryon resonances
The Statistical Hadronization Model (SHM) assumes that hadrons are produced according to the density of thermal yields $N_{h}^{th}(T)$. In ultra-relativistic heavy-ion collisions, the SHM successfully reproduces the yields of meson and baryon ground states with one chemical freeze-out temperature of $T_{cf}\simeq 156$ MeV [67], which represents the temperature of the medium at which the different particle abundances get fixed. The other model parameters are the baryo-chemical potential $\mu_{B}$ and the volume of the systems, which are found to be $\mu_{B}$ = 0.7 $\pm$ 3.8 MeV, and V = 5280 $\pm$ 410 $\mathrm{f}m^{3}$. Its extension to charm (and beauty) hadrons needs to consider that the charm quarks are produced in the initial stage of the collisions with a multiplicity that is quite larger than the one expected from thermal charm production at $T_{cf}$. To this purpose, a charm-quark fugacity factor, $g_{\mathrm{c}}$, is introduced in the model to account for the excess of the initial charm yield wrt to the thermal one. A charm balance equation is needed to guarantee the conservation of charm quarks for a given value of $g_{c}$, while allowing the relative abundances of the different open and hidden charm hadrons to follow the statistical thermal ones, it can be written as:
@@
where $N_{\mathrm{c}\bar{c}}$ in the total charm yield in the rapidity range considered, $n_{i,j,k}^{{\mathrm{t}h}}$ are the (grand-canonical) thermal yields for open charm with one charm quark ($h_{oc,1}$) and two charm quarks ($h_{oc,2}$) and for hidden charm ($h_{hc}$). In principle, Eq. 5 should include the hadrons with three charm quarks but, in the current implementations, they are discarded under the reasonable assumption that their contribution to the sum is negligible. Also, from Eq. (5) we can see that the fugacity $g_{c}$ can be seen as proportional to the ratio of total charm yield to the thermal charm one, in the limit where multicharm and hidden charm hadron yields are negligible. Fugacity factors in elementary collisions are introduced to account for the non-equilibrium yield of quarks heavier than $u$ and $d$, thus for strangeness as well as for charm. However, due to the dominance of hadrons with one charm quark this does not affect baryon-to-meson yield ratios of single charm hadrons, because both baryons and mesons contain one charm. On the other hand, one can also notice that doubling $N_{\mathrm{c}\bar{c}}$ implies doubling $g_{\mathrm{c}}$, with a consequent enhancement by a factor of 2 of single-charm hadrons and a factor of 4 increase of double-charm hadrons. Such a scaling with $N_{\mathrm{c}\bar{c}}$ is expected also in the coalescence model described in the next paragraph.
@@ II.2.4 Hadronization via coalescence
with $g_{H}$ indicating the statistical factor giving the probability that two (or three) random quarks have the right quantum numbers to match the quantum number of the hadron (flavour, spin, isospin) in a colourless combination. For D mesons the statistical factor $g_{\mathrm{D}}=1/36$ gives the probability that two random quarks have the right colour, spin, and isospin to match the quantum number of the considered mesons. For $\Lambda_{\mathrm{c}}^{+}$ the statistical factor is $g_{\Lambda_{\mathrm{c}}^{+}}=1/108$. $N_{\mathrm{q}}$ is the number of quarks that form the hadron, $N_{\mathrm{q}}=2$ for mesons and $N_{\mathrm{q}}=3$ for baryons. The $d\sigma_{i}$ denotes an element of a space-like hypersurface, while $f_{{\mathrm{q}}_{i}}$ are the quark (anti-quark) phase-space distribution functions for $i^{\mathrm{th}}$ quark (anti-quark) and the coalescence function ${\cal{C}}_{H}=C^{N_{\mathrm{q}}-1}f_{H}(x_{1}...x_{N_{{\mathrm{q}}}},p_{1}...p_{N_{{\mathrm{q}}}})$ is directly proportional to the Wigner function which describes the spatial and momentum distribution of quarks in a hadron, $f_{H}$. The constant $C^{N_{\mathrm{q}}-1}$ is an important and specific aspect of the Catania model, not present in the standard coalescence approach. It is this constant that allows normalizing the coalescence probability for charm quarks such that the total coalescence probability (i.e. summing up all the mesons and baryons channel considered) in the limit $p_{\mathrm{T}}\rightarrow 0$ is $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$. Such a constant $C$ can be determined only by calculating all the charmed mesons and baryons. In the first work [55] it was evaluated considering only D, $\mathrm{D^{*+}}$, $\mathrm{D_{s}^{+}}$, $\Lambda_{\mathrm{c}}^{+}$, and $\Sigma_{\mathrm{c}}^{0,+,++}$ hadrons. However, more recently the production of $\Xi_{\mathrm{c}}^{+,0}$ and $\Omega_{\mathrm{c}}^{0}$ was observed not to be negligible [54]. Having fixed such a normalization, the probability of a charm quark to hadronize via coalescence or fragmentation depends as in the standard approach on its transverse momentum according to Eq. 6. The coalescence probability is high for low quark momenta and quickly decreases with increasing transverse momentum, meaning that most of charm quarks hadronize via coalescence at low $p_{\mathrm{T}}$ whereas, in the high momentum region, fragmentation is the dominant contribution. Imposing the condition $P_{coal}(p\rightarrow 0)=1$ is a key aspect underlying the prediction of quite large values of heavy baryons: $\Lambda_{\mathrm{c}}^{+}/\mathrm{D^{0}}\sim O(1)$, $\Xi_{\mathrm{c}}^{0}/\mathrm{D^{0}}\sim 0.2$ and $\Omega_{\mathrm{c}}^{0}/\mathrm{D^{0}}\sim 0.1$ [54]. The physical motivation is related to the fact that in the standard coalescence approach there is no dynamics related to confinement. The renormalization constant $C$ appears linearly for mesons and quadratic for baryons. This is physically motivated at least by two considerations. First, enforcing $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ implies that there is an enhancement of the recombining probability (with respect to the one from a simple phase space overlap of the wave functions) that can be associated to the dynamics of the confining interaction that increases the coalescence probability (here accounted for just by a renormalization of a coupling constant). Such an effect has to be squared for baryons for which two relative distances among the quarks have to be defined. Second, one has to consider that the density of the medium at which the hadrons are formed is not known, and being the phase transition a cross-over it is likely that there is a density interval over which it occurs. Therefore the hadrons yields and ratios in a standard coalescence approach (i.e. without imposing $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$) would significantly depend on the specific value of the density chosen. In the Catania approach the probability of forming a hadron scales linearly with the density for mesons and quadratically for baryons. If one chooses a lower density value, implying a reduced standard coalescence probability, the constraint $P_{coal}(p_{\mathrm{T}}\rightarrow 0)=1$ increases the value of the coefficient $C^{N_{q}-1}$, making the yields and the baryon-over-meson ratios quite stable and nearly independent of the specific value of the density of the background fireball employed to perform the calculation.
@@
with $x_{ri}$ and $p_{ri}$ relative coordinates in space and momentum. We notice that an important aspect of a proper implementation of the coalescence approach is that, respecting the basic feature of quantum mechanics, the width $\sigma_{r}$ of the wave function in r-space constrains also the Wigner function width in p-space; for a Gaussian wave function for each relative coordinate we have $\sigma_{ri}\cdot\sigma_{pi}=1$ and $\sigma_{p}$ cannot be used as a further independent parameter. It can be noticed that the coalescence process, and the related formula, can be factorized as a diquark recombination followed by a diquark-quark coalescence if the diquark binding energy is vanishing.
It is interesting to discuss the dependence of the coalescence results on the specific value of the wave-function width, which is the main parameter of the model. This can be easily evaluated analytically considering a non-relativistic reduction, which, even if not quantitatively exact, allows having a semi-quantitative estimate of the main aspect involved in the coalescence process. Assuming a relativistic boost invariant thermal distribution function homogeneous in $r$-space, considering the mid-rapidity region and employing a non-relativistic reduction for the quark energy $E=m+p^{2}/2m$, discarding the radial flow, and assuming a Gaussian wave function, it can be shown that the coalescence yield for mesons and baryons scales as: